The investor is presented with the two following stocks:

Expected Return

Standard Deviation

Stock A

10%

30%

Stock B

20%

60%

What is the expected return on the portfolio that invests 30% in stock A?

A.

15%

B.

17%

C.

27%

D.

23%
Answer #1

Weight of Stock B=(100-30)=70%

Expected return=Respective return*Respective weight

=(0.3*10)+(0.7*20)

which is equal to

=17%